### v1

# DReSS: Data-driven Regularized Structured Streamlining for Large Language Models

Abstract
Large language models (LLMs) have achieved significant progress across various domains, but their increasing scale results in high computational and memory costs. Recent studies have revealed that LLMs exhibit sparsity, providing the potential to reduce model size through pruning techniques. However, existing pruning methods typically follow a prune-then-finetune paradigm. Since the pruned components still contain valuable information, their direct removal often leads to irreversible performance degradation, imposing a substantial computational burden to recover performance during finetuning. In this paper, we propose a novel paradigm that first applies regularization, then prunes, and finally finetunes. Based on this paradigm, we introduce DReSS, a simple and effective Data-driven Regularized Structured Streamlining method for LLMs. By leveraging a small amount of data to regularize the components to be pruned, DReSS explicitly transfers the important information to the remaining parts of the model in advance. Compared to direct pruning, this can reduce the information loss caused by parameter removal, thereby enhancing its language modeling capabilities. Experimental results demonstrate that DReSS significantly outperforms existing pruning methods even under extreme pruning ratios, significantly reducing latency and increasing throughput.
Machine Learning, ICML
1 Introduction
Large language models (LLMs) have achieved significant advancements across a wide range of tasks and domains, demonstrating their robust capabilities (Zhang et al., 2022; Achiam et al., 2023; Touvron et al., 2023; Wu et al., 2024). However as the model size grows, the increased number of parameters results in substantial computational and memory demands. These factors severely hinder the deployment of LLMs in practical applications. Therefore, it is urgent to develop methods that can reduce model size while maintaining performance.
To address these challenges, several methods have been proposed, including pruning (Frantar & Alistarh, 2023; Sun et al., 2023; Ma et al., 2023; An et al., 2024), quantization (Frantar et al., 2022; Xiao et al., 2023), knowledge distillation (Shridhar et al., 2022; Hsieh et al., 2023), and low-rank decomposition (Saha et al., 2023). In this work, we mainly focus on pruning, an efficient and highly generalizable method that can be combined with other model compression strategies. Pruning techniques are generally classified into two primary categories: unstructured pruning (Frantar & Alistarh, 2023; Sun et al., 2023) and structured pruning (Ma et al., 2023; An et al., 2024). Compared to unstructured pruning, structured pruning offers the flexibility to do recovery fine-tuning (RFT) on the pruned model for specific downstream tasks without relying on specialized hardware (Zhu et al., 2024). Moreover, the model obtained through structured pruning typically achieves much faster inference speed, while ensuring the same perplexity and accuracy as the model obtained through unstructured pruning.
Despite these advancements, existing structured pruning methods still have some limitations. They all follow the paradigm of first selecting channels or layers for pruning based on a designed metric, and then performing RFT (Chavan et al., 2024). However, they neglect the fact that important information can also exist in the pruned channels or layers (Dettmers et al., 2022; Xiao et al., 2023; Yin et al., 2023), and directly removing them leads to an irreversible decline in model performance. Thus, in order to achieve satisfactory performance with the pruned model, a large amount of data must be introduced for RFT (Ma et al., 2023), which increases computational overhead. Moreover, they often lead to a collapse in model performance at high pruning ratios, limiting their effectiveness in reducing lantency and enhancing throughput.
To tackle these limitations, in this paper, we propose a new pruning paradigm that first applies regularization to the components to be pruned, then removes those parameters, and finally performs RFT. To the best of our knowledge, we are the first to propose this paradigm. The regularization process explicitly transfers important information from the parts to be pruned to the remaining parts, thereby reducing the information loss caused by direct parameter removal. Based on this paradigm, we introduce DReSS, a Data-driven Regularized Structured Streamlining method for LLMs. The comparison of the proposed method with existing channel-wise pruning approaches is shown in Figure 1. Specifically, the process of DReSS is divided into the following steps: (1) select data for pre-pruning regularization and post-pruning RFT, (2) apply regularization to the selected channels in the parameter matrix, and (3) prune the specified channels, then (4) do RFT. First, we randomly selecting a small subset of data from widely used benchmark datasets, as the amount of selected data is small, the overhead in the regularization and RFT process can be significantly reduced. Second, we determine the channels to be pruned based on the pruning ratio, and apply regularization to these channels using the selected data. This shifts important information to the unpruned parts, significantly mitigating performance degradation caused by direct pruning and ensuring the model maintains strong performance even at high pruning ratios, thus enabling substantial inference speedup. Third, we prune the parameters that have been regularized. Finally, we do RFT on the pruned model using a small subset of samples selected in the first step. Extensive experiments demonstrate that, DReSS significantly surpasses existing pruning methods in terms of perplexity and accuracy across various pruning ratios, it also substantially improves inference speed. The main contributions of our work are summarized as follows:
  • Propose A New Paradigm: By first applying regularization, then pruning, and finally performing RFT, DReSS reduces information loss from direct parameter removal, enhancing the model’s language modeling ability and lowering the computational cost of RFT.
  • High Performance: DReSS outperforms existing structured pruning methods on both generation and zero-shot tasks. The pruned model shows a significant increase in throughput and a notable reduction in latency compared to the dense model.
  • Low Overhead: By using only a small amount of data during the regularization and optional RFT process, the pruning overhead is minimal.
Figure 1: A comparison between existing channel-wise structured pruning methods and DReSS. The deeper blue square represents greater performance impact; the red exclamation mark indicates the important information. The taller orange-yellow cylinder represents larger data volume. Here, L($\cdot$) represents the regularization, and S represents pruning. Above: Existing methods first select channels based on importance, followed by pruning and RFT on the pruned model. Below: DReSS first regularizes channels to minimize performance impact, prunes the channels, then do RFT with a little amount of data.
2 Related Works
2.1 Pruning Methods
The elimination of redundant weights, referred to as pruning, has been an effective approach for reducing the complexity of deep neural networks over the past several decades (LeCun et al., 1989; Hassibi et al., 1993; Han et al., 2015; Wen et al., 2016; Louizos et al., 2017). Pruning methods can be broadly categorized into two types: unstructured pruning (Kurtic et al., 2022; Zhang et al., 2024; Xu et al., 2024) and structured pruning (Xia et al., 2023b; Gao et al., 2024b; Yang et al., 2024).
Unstructured pruning methods remove individual weights based on their importance. For instance, Magnitude pruning (Han et al., 2015; Zhou et al., 2021) assumes larger weights are more important, Wanda (Sun et al., 2023) uses the product of layer weights and activations as the metric, and SparseGPT (Frantar & Alistarh, 2023) relies on the weight squared divided by the inverse Hessian matrix. The main advantage is flexibility, enabling the removal of unimportant weights while retaining critical ones, allowing high sparsity in LLMs. However, its limitations include the need for specialized hardware for accelerated inference (Xia et al., 2023a), high sparsity for significant acceleration (Wang, 2020), and inability to perform RFT on downstream tasks.
Structured pruning methods, including channel-wise pruning (An et al., 2024) and layer-wise pruning (Men et al., 2024), can achieve significant inference speedup with a low sparsity. Channel-wise pruning methods typically design a metric to evaluate the importance of channels in the parameter matrix of LLMs and then prune the less important channels. For instance, SliceGPT (Ashkboos et al., 2024) adopts Principal Component Analysis (PCA), using the eigenvalues of the covariance matrix of activation values as a metric to preserve channels corresponding to larger eigenvalues. LLM Surgeon (van der Ouderaa et al., 2023) periodically updates model weights and structures, pruning more aggressively in the initial layers and less so in the intermediate layers. Layer-wise pruning methods, such as SLEB (Song et al., 2024), iteratively prune entire transformer layers by assessing their importance. However, these methods have a key limitation: even less important channels or layers may contain valuable information, and pruning them directly often leads to significant performance degradation. To address this, we investigate the possibility of ‘shifting’ important information from the pruned parts to the retained parts, which could substantially improve model performance.
2.2 Regularization
Regularization is widely employed in machine learning (Hoerl & Kennard, 1970; Poggio et al., 1987; Balestriero et al., 2022), such as feature selection (Tibshirani, 1996) and preventing model overfitting (Santos & Papa, 2022). The $\ell_{1}$-norm typically induces sparsity by driving certain coefficients to zero, while $\ell_{2}$-norm encourages smoother solutions (Boyd & Vandenberghe, 2004). Both of them can significantly alter the distribution pattern of the data (Han et al., 2015; Tao et al., 2023). Motivated by this observation, we propose a data-driven regularization approach for parameters designated for pruning. By applying a small amount of data, the regularization probably transfers important information from the pruned parameter space to the retained parts of the model, thereby enhancing the performance of the pruned model.
3 Methodology
In this section, we provide a detailed introduction to DRess. The overall DReSS algorithm is summarized in Algorithm 1, our method consists of four main steps:
  • Data Selection: Select a small amount of data for pre-pruning regularization and post-pruning RFT.
  • Regularization: Applying regularization to the selected channels of weight matrices during both forward and backward propagations.
  • Pruing: Prune the regularized portions.
  • RFT: Perform RFT on the pruned model using the subset of data selected in the first step.
Algorithm 1 DReSS algorithm.
  Input: selected data $X$, number of layers in LLMs $l$, initial model $W$, number of regularization iterations $n$, norm type: flag.
  if flag==1 then
     use Proposition 3.2 to transform problem
  end if
  $X_{1},X_{2}\leftarrow divide(X)$
  for $i=1$ to $n$ do
[…]
     ~$\tilde{\mathcal{L}}_{\text{att}}$$\leftarrow$00, ~$\tilde{\mathcal{L}}_{\text{FFN}}$$\leftarrow$00, ~$\tilde{\mathcal{L}}_{\text{2}}$$\leftarrow$00
     for $j=1$ to $l$ do
        ~$\tilde{\mathcal{L}}_{\text{att}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{att}}$+$regularize(W_{Q}^{j},W_{K}^{j},W_{V}^{j},W_{o}^{j},X_{1})$
        ~$\tilde{\mathcal{L}}_{\text{FFN}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{FFN}}$+$regularize(W_{up}^{j},W_{down}^{j},X_{1})$
        ~$\tilde{\mathcal{L}}_{\text{2}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{2}}$+$regularize(w_{N}^{j},X_{1})$
     end for
[…]
     ~$\tilde{\mathcal{L}}_{\text{remain}}$$\leftarrow$~$\tilde{\mathcal{L}}_{\text{1}}$+~$\tilde{\mathcal{L}}_{\text{2}}$+~$\tilde{\mathcal{L}}_{\text{3}}$
     $\mathcal{L}(W,X)\leftarrow forward(W,X_{1})$
     $\mathcal{L}_{\text{sum}}$$\leftarrow$$\mathcal{L}(W,X)$+~$\tilde{\mathcal{L}}_{\text{att}}$+~$\tilde{\mathcal{L}}_{\text{FFN}}$+~$\tilde{\mathcal{L}}_{\text{remain}}$
     update $W$ using BP algorithm
  end for
  $W\leftarrow Prune(W,R)$
  $W\leftarrow Optional\_RFT(W,X_{2})$
3.1 Data Selection
The required data is selected from widely used calibration datasets, including Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020). For instance, approximately 1,000 samples are randomly selected from the WikiText-2 training dataset for both pre-pruning regularization and post-pruning RFT. In Section 4.5, we analyze the impact of different datasets.
3.2 Regularization
To provide a clearer understanding of this section, we first introduce the relevant notations. Let $p$ denote the pruning ratio, which indicates that $p\%$ of the model’s parameters need to be pruned. The model dimension of the LLMs is represented by $d$, and the number of layers is denoted by $l$. Additionally, let $W_{emb}\in\mathbb{R}^{b\times n\times d}$ represents the data feature map, and $W_{pos}\in\mathbb{R}^{b\times n\times d}$ denotes the positional feature map, where $b$ is the batch size and $n$ is the number of tokens. The model weights are represented by $W\in\mathbb{R}^{d_{1}\times d_{2}}$. Specifically, $W_{Q}^{i}\in\mathbb{R}^{d\times d_{1}}$, $W_{K}^{i}\in\mathbb{R}^{d\times d_{1}}$, $W_{V}^{i}\in\mathbb{R}^{d\times d_{1}}$, and $W_{o}^{i}\in\mathbb{R}^{d_{1}\times d}$ represent the parameter matrices in the $i$th layer Attention block. Similarly, $W_{up}^{i}\in\mathbb{R}^{d\times d_{2}}$ and $W_{down}^{i}\in\mathbb{R}^{d_{2}\times d}$ represent the parameter matrices in the $i$th layer FFN block. Additionally, $w_{N}^{i}\in\mathbb{R}^{d}$ denotes the LayerNorm vector in the $i$-th layer, and $W_{lm}\in\mathbb{R}^{b\times n}$ represents the language modeling matrix. Finally, $I\in\mathbb{R}^{d\times d}$ represents the identity matrix, let $R\in\mathbb{R}^{d\times d}$ represent a pseudo-index selection matrix, which is a diagonal matrix consisting of $(1-p\%)\times d$ zeros and $p\%\times d$ ones, where the ones correspond to the selected indices.
Figure 2: The pruning strategy of DReSS in a transformer layer.
Proposition 3.1.
(Dependency between the Attention Block and FFN Block.) If regularization is applied to certain columns of $W_{down}^{i-1}$ in the $(i-1)$th layer FFN block, then regularization must also be applied to the corresponding rows of $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ in the $i$th layer Attention block. The same dependency applies to $W_{o}^{i}$ and $W_{up}^{i-1}$.
We illustrate the process of applying regularization within a transformer layer (Vaswani, 2017), as shown in Figure 2. Following Proposition 3.1, in the Attention block, regularization is applied to the rows of $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ corresponding to the one entries in the index matrix $R$, as well as to the columns of $W_{o}^{i}$ associated with these entries. Similarly, in the FFN block, regularization is enforced on the rows of $W_{up}^{i}$ and the columns of $W_{down}^{i}$ that correspond to the one entries in $R$. We also apply regularization to the remaining parts in LLMs. Specifically, for $W_{emb}$, regularization is enforced on the columns corresponding to the one indices in $R$, and the same procedure is applied to $W_{pos}$. For each $w_{N}^{i}$ within every transformer layer, regularization is applied to the elements corresponding to the one indices in $R$. Additionally, for the final $W_{lm}$, due to its dependency on $W_{down}^{l}$ in the FFN block of the last transformer layer, regularization is applied to the corresponding rows. Therefore, the overall loss function consists of the language modeling loss and the loss induced by regularization, which is defined below:
$\begin{split}\mathcal{L}_{\text{sum}}=\mathcal{L}(W,X)+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{att}}^{i}+\sum_{i=1}^{l}\tilde{\mathcal{L}}_{\text{FFN}}^{i}+\tilde{\mathcal{L}}_{\text{remain}}\end{split}$ (1)
where $\mathcal{L}(W,X)$ denotes the language modeling loss and $X$ is the input data, and the remaining term represents the regularization loss. ~$\tilde{\mathcal{L}}_{i}^{\text{att}}$ denotes the $i$th layer attention block regularization loss, which is defined as:
~$\begin{split}\tilde{\mathcal{L}}_{i}^{\text{att}}=\lambda(\|RW_{Q}^{i}\|+\|RW_{K}^{i}\|+\|RW_{V}^{i}\|+\|W_{o}^{i}R\|)\end{split}$ (2)
~$\tilde{\mathcal{L}}_{i}^{\text{FFN}}$ denotes the $i$th layer FFN block regularization loss and the definition is as below:
~$\begin{split}\tilde{\mathcal{L}}_{\text{FFN}}^{i}=\lambda\|RW_{up}^{i}\|+\lambda\|W_{down}^{i}R\|\end{split}$ (3)
~$\tilde{\mathcal{L}}_{\text{remain}}$ denotes the remaining parts of the LLMs’ regularization loss as follows:
~$\begin{split}\tilde{\mathcal{L}}_{\text{remain}}=\lambda(\|W_{emb}R\|+\|W_{pos}R\|+\\ \sum_{i=1}^{l}\|Rw_{N}^{i}\|+\|RW_{lm}\|)\end{split}$ (4)
[…]
Proposition 3.2.
(If the loss function includes the $\ell_{1}$-norm, it can also be solved using backpropagation.) The following unconstrained optimization problem is equivalent to the constrained optimization problem, where $|\cdot|_{1}$ denotes the $\ell_{1}$-norm.
$\min\quad|x|_{1}\quad\iff\quad\min_{x,y}$ $\quad\mathbf{1}^{T}y$ (5)
s.t. $\quad-y\leq x\leq y,$
$\quad y\geq 0.$
We formalize the minimization of the overall loss as an optimization problem. When $\ell_{2}$-norm is used, the problem can be efficiently solved using backpropagation. In contrast, when the $\ell_{1}$-norm is employed, as the regularization losses introduced by various components in LLMs are similar, we consider only ~$\tilde{\mathcal{L}}_{\text{FFN}}^{i}$ as an example. According to Proposition 3.2, the optimization problem in Equation 6 can be equivalently transformed into a constrained formulation in Equation 7, enabling its solution by backpropagation.
$\min_{W}\quad\mathcal{L}(W,X)+\lambda|RW_{up}^{i}|_{1}+\lambda|W_{down}^{i}R|_{1}$ (6)
$\min_{W,Y_{1},Y_{2}}$ $\quad\mathcal{L}(W,X)+\mathbf{1}^{T}Y_{1}\mathbf{1}+\mathbf{1}^{T}Y_{2}\mathbf{1}$ (7)
s.t. $\quad-Y_{1}\leq RW_{up}^{i}\leq Y_{1},$
$\quad-Y_{2}\leq W_{down}^{i}R\leq Y_{2},$
$\quad Y_{1}\geq 0,\quad Y_{2}\geq 0.$
During the forward and backward propagation processes, due to the application of regularization on certain specific channels in LLMs, the values within these rows or columns are significantly reduced. Intuitively, this reduction could potentially decrease the impact of these parameters on the model’s performance.
3.3 Pruning
After completing the regularization, we prune the rows and columns directly using the pseudo-index selection matrix $R$. For clarity, we define $S=I-R$ and the pruning for the $i$th layer attention block is then performed as:
${W_{Q}^{i}}^{\prime}$ $=SW_{Q}^{i},\quad{W_{K}^{i}}^{\prime}=SW_{K}^{i}$ (8)
${W_{V}^{i}}^{\prime}$ $=SW_{V}^{i},\quad{W_{o}^{i}}^{\prime}=W_{o}^{i}S$
While for the FFN block, it follows:
${W_{up}^{i}}^{\prime}=SW_{up}^{i},\quad{W_{down}^{i}}^{\prime}=W_{down}^{i}S$ (9)
The remaining parts are pruned as below:
${W_{emb}}^{\prime}$ $=W_{emb}S,\quad{W_{pos}}^{\prime}=W_{pos}S$ (10)
${w_{N}^{i}}^{\prime}$ $=S{w_{N}^{i}},\quad{W_{lm}}^{\prime}=SW_{lm}$
The entire pruning process is straightforward and simple.
3.4 Optional RFT
After pruning, we perform RFT on the model using a small subset of the data selected in Section 3.1, leveraging LoRA (Hu et al., 2021). This step is optional, as experimental results in section 4.7 show that DReSS without RFT still maintains competitive performance. in terms of both perplexity and zero-shot accuracy.
Table 1: Performance comparison of different pruning methods. ’PPL’ refers to the perplexity on WikiText-2. The accuracy is reported on five zero-shot tasks: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c, as well as the average accuracy. The best result is highlighted in bold, and the second-best in underlined. DReSS-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, while DReSS-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm.
Model Method Pruning Ratio PPL ($\downarrow$) PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc Phi-2 Dense 0% 5.28 79.11 75.77 73.83 78.32 54.18 72.24 LLM Surgeon 25% 7.26 67.28 63.25 54.24 51.62 35.85 54.44 SliceGPT 25% 7.08 69.21 65.35 52.40 53.70 31.66 54.46 SLEB 25% 7.82 67.94 62.79 49.80 48.55 29.34 51.68 DReSS-$\ell_{2}$ 25% 6.25 68.52 66.71 56.73 52.78 37.63 56.47 DReSS-$\ell_{1}$ 25% 6.28 68.67 65.32 57.16 52.39 37.28 56.16 OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 LLM Surgeon 25% 12.60 68.22 58.06 46.25 51.52 27.31 50.27 SliceGPT 25% 14.56 65.29 57.22 47.85 49.79 27.99 49.63 SLEB 25% 16.73 63.28 55.37 43.09 45.62 24.21 46.31 DReSS-$\ell_{2}$ 25% 12.52 69.12 58.98 49.10 50.62 28.15 51.20 DReSS-$\ell_{1}$ 25% 12.48 68.89 58.33 48.62 51.67 27.55 51.01 OPT-6.7B Dense 0% 10.85 76.39 65.19 67.16 60.14 34.64 60.70 LLM Surgeon 25% 11.33 72.12 61.08 62.36 54.19 30.52 56.05 SliceGPT 25% 11.90 70.35 60.62 58.15 52.78 29.52 54.28 SLEB 25% 13.80 71.28 60.96 59.26 51.54 29.83 54.57 DReSS-$\ell_{2}$ 25% 11.06 74.26 63.12 61.26 56.64 31.77 57.41 DReSS-$\ell_{1}$ 25% 11.74 73.86 64.28 60.45 56.36 32.07 57.40 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 LLM Surgeon 25% 11.02 74.18 64.33 65.37 60.96 34.98 59.96 SliceGPT 25% 10.94 73.67 64.25 63.28 60.52 34.64 59.27 SLEB 25% 12.02 72.62 63.96 62.79 59.21 34.12 58.50 DReSS-$\ell_{2}$ 25% 10.38 74.06 64.57 66.82 61.56 35.33 60.47 DReSS-$\ell_{1}$ 25% 10.56 73.65 64.38 66.46 61.15 34.74 60.07 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 LLM Surgeon 25% 7.38 70.59 65.87 58.66 63.65 38.33 59.42 SliceGPT 25% 7.55 66.87 63.38 54.16 58.46 34.56 55.48 SLEB 25% 10.24 63.25 62.36 53.77 55.82 32.24 53.49 DReSS-$\ell_{2}$ 25% 5.86 73.18 66.49 61.73 65.42 40.86 61.54 DReSS-$\ell_{1}$ 25% 5.81 73.42 65.73 61.92 65.26 39.68 61.20 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 LLM Surgeon 25% 5.75 77.75 69.62 74.31 72.83 43.52 67.61 SliceGPT 25% 6.30 68.55 67.48 58.10 62.50 37.88 58.90 SLEB 25% 7.39 66.93 65.87 55.48 60.06 35.14 56.70 DReSS-$\ell_{2}$ 25% 5.12 76.14 71.22 76.51 73.45 46.87 68.84 DReSS-$\ell_{1}$ 25% 5.16 77.24 70.61 76.75 73.84 45.46 68.78
Figure 3: The ratio of the sum of absolute parameter values after regularization to that before regularization for LLaMA2-7B. For example, regularization is applied to the last 25% of rows in $W_{Q}$, and the absolute sum of these rows after regularization is divided by the sum of the corresponding rows before regularization to obtain the Preserved Ratios. This is applied similarly to other matrices such as $W_{K}$, $W_{V}$, $W_{o}$, $W_{gate}$, $W_{up}$, $W_{down}$, and for layer activations $X$, the sum of the regularized columns is compared to the original sum.
4 Experiments
This section introduce experimental setup 4.1 and analyzes the effectiveness of DReSS from the following aspects: performance 4.2, acceleration effectiveness 4.3, robustness under different pruning ratios 4.4, robustness under different datasets 4.5, low overhead 4.6, and ablation study 4.7.
4.1 Experimental Setup
Implementation DReSS and baseline methods are implemented on the PyTorch framework (Paszke et al., 2019), utilizing the Hugging Face Transformers library (Wolf, 2019). All experiments are conducted on 80GB NVIDIA A100 GPUs. For fairness, we use llm-eval-harness (Gao et al., 2024a) to evaluate the pruned models.
Datasets To evaluate the performance of DReSS, we conducted experiments on generation and zero-shot tasks. For generation task, following prior work(Ashkboos et al., 2024), we evaluate the model’s perplexity on WikiText-2. For zero-shot task evaluation, the benchmarks consists of PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), HellaSwag (Zellers et al., 2019), ARC-e and ARC-c (Clark et al., 2018). To demonstrate that DReSS does not have a specific dependency on the regularization data, we use a random selection of data from Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020) for calibration prior to pruning and for the RFT after pruning.
Models The models pruned using DReSS include the LLaMA models (LLaMA2-7B, LLaMA2-13B) (Touvron et al., 2023), OPT models (OPT-2.7B, OPT-6.7B, OPT-13B) (Zhang et al., 2022), and Phi-2 (Javaheripi et al., 2023).
Baselines We evaluate DReSS against structured pruning methods: LLM Surgeon (van der Ouderaa et al., 2023), SliceGPT (Ashkboos et al., 2024), and SLEB (Song et al., 2024). Among them, LLM Surgeon and SliceGPT are channel-wise, SLEB is layer-wise. To ensure a fair comparison, in the subsequent experiments, we ensure that the data used for regularization in DReSS is consistent with the calibration data used by LLM Surgeon, SliceGPT, and SLEB before pruning, and that the data used for RFT in each method is also the same.
Evaluation Metrics The performance on generation task is measured by perplexity, a well-established and robust metric (Yao et al., 2022), while zero-shot tasks performance is evaluated using accuracy (Dong et al., 2024). The acceleration effects are represented by throughput and latency (Song et al., 2024).
4.2 Performance Comparison
To ensure fairness, we randomly selected 1,000 samples from WikiText-2 for each method with a sequence length of 2048. The ratio of calibartion data and the RFT data was set to 3:1, and the pruning ratio was 25%.
As shown in Table 1, DReSS demonstrates superior performance across both generation and zero-shot tasks achieving lowest perplexity and highest average accuracy across all models. For instance, on LLaMA2-7B, its perplexity is 20% lower than the second-best method, LLM Surgeon. Notably, on OPT-13B, compared to the dense model, DReSS experiences only a 1% drop in average accuracy. The performance difference between the $\ell_{2}$-norm and $\ell_{1}$-norm is tiny. In Figure 3, we present the ratio of the absolute values of model parameters and activation values after and before regularization of the selected portions. The results show that, in most intermediate layers, the parameters and activation values after regularization are reduced to 30% of their original values, which indicates that the amount of information contained in this part of the parameter space has decreased. Overall, regularization may help transfer important information hidden in the pruned portions to the remaining parts.
4.3 Acceleration Comparison
Language processing in LLMs comprises two primary stages: prompt processing, which is compute-bound, and token generation, which is memory-bound. We separately analyze the speedup achieved in each stage. Table 2 presents the throughput and latency results for OPT-13B and LLaMA2-13B, evaluated using a single NVIDIA A100 GPU. Following the methodology of previous work (Song et al., 2024), the token generation test scenario involves generating sentences of 128 tokens with a batch size of 64, whereas for prompt processing, latency is measured by processing an input sequence of 2048 tokens.
At a pruning ratio of 50% on OPT-13B, DReSS delivers a 35% improvement in throughput and a 30% reduction in latency compared to the dense model. These results highlight the superior efficiency of DReSS in accelerating model inference.
Table 2: Comparison of throughput and latency under different ratios on OPT-13B and LLaMA2-13B. ‘PPL’ refers to the perplexity on Wikitext2, ‘PR’ represents ‘pruning ratio’, ‘TI’ represents ‘throughput increase’.
Model Method PR PPL($\downarrow$) Tokens/s TI Latency(ms) Speedup OPT-13B Dense 0% 10.12 1029 1.00× 386.5 1.00× DReSS 25% 10.38 1194 1.16× 319.42 1.21× DReSS 50% 12.15 1389 1.35× 274.11 1.41× LLaMA2-13B Dense 0% 4.88 1066 1.00× 396.9 1.00× DReSS 25% 5.12 1215 1.14× 330.8 1.20× DReSS 50% 7.59 1407 1.32× 285.5 1.39×
4.4 Robustness to Varying Pruning Ratios.
Keeping all other settings consistent with Section 4.2, we extend the pruning ratio from 20% to 60%. The perplexity of LLaMA2-7B on WikiText-2 under different methods are shown in Figure 4. DReSS significantly outperforms other methods across various pruning ratios. At a pruning ratio of 20%, DReSS performs similarly to the dense model. When the pruning ratio is up to 60%, SLEB collapses, while DReSS maintains relatively low perplexity compared to other pruning methods. This demonstrates that DReSS maintains robust performance even under extreme pruning ratios, enabling structured pruning to achieve high sparsity levels and unlocking significant potential for model acceleration.
Figure 4: Perplexity of LLaMA-7B pruned by various approaches under different pruning ratios on WikiText-2.
4.5 Dependency on Datasets
Since DReSS relies on data-driven regularization, we investigate its potential dataset dependence. We evaluated perplexity for four methods on WikiText-2, using calibration and RFT data selected from Alpaca, WikiText-2, PTB, and C4. To ensure fairness, we randomly selected 1,000 samples from each dataset, with other settings consistent with Section 4.2. As shown in Figure 5, DReSS consistently outperforms the other methods across datasets, demonstrating its robustness and scalability.
Figure 5: Comparison of perplexity on different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.
4.6 Minimal Overhead
We evaluated the perplexity of each method on LLaMA2-7B using varying amounts of data, keeping all other conditions consistent with Section 4.2 and only change the data size from 500 to 8,000. As shown in Table 3, when only 1,000 samples were used for regularization and RFT, DReSS outperformed the other methods that used 4,000 samples. This further highlights the effectiveness and efficiency of applying regularization prior to pruning, as it transfers critical information from the pruned segments in advance, enabling DReSS to deliver strong performance with minimal data overhead.
Table 3: The impact of the amount of data on model performance.
Data Size 500 1000 2000 4000 8000 LLM Surgeon 8.14 7.38 6.55 5.89 5.56 SliceGPT 8.63 7.55 6.82 6.07 5.74 SLEB 13.26 10.24 9.67 8.79 7.83 DReSS 6.75 5.86 5.61 5.48 5.43
4.7 Ablation Study
Effectiveness of Regularization As shown in Table 4, both random pruning and full-parameter regularization significantly increase perplexity on WikiText-2 and reduce average accuracy at a 25% pruning ratio for LLaMA2-7B and LLaMA2-13B. This underscores the importance of applying regularization only to the pruned components. Additionally, RFT after pruning has minimal impact on perplexity and average accuracy, indicating that the RFT process is optional.
Table 4: Ablation results on LLaMA2-7B and LLaMA2-13B. ‘FPR’ refers to full parameter regularization, ‘R’ refers regularization. “w/o” denotes variants with specific components removed.
Model Method Setting PPL($\downarrow$) $\Delta$ AVG_ACC $\Delta$($\downarrow$) LLaMA2-7B DReSS 5.86 0.00 61.54 0.00 w FPR 12.68 +6.82 50.26 -11.28 w/o R 22.38 +16.52 47.25 -14.29 w/o RFT 5.97 + 0.11 59.56 -1.98 LLaMA2-13B DReSS 5.12 0.00 68.84 0.00 w FPR 10.53 +5.41 53.28 -15.56 w/o R 18.94 +13.82 50.17 -18.67 w/o RFT 5.39 + 0.27 67.96 -0.88
Table 5: The perplexity on WikiText-2 under different regularization and RFT data ratio using LLaMA2-7B at a pruning ratio of 25%.
Ratio (R/F) 4:1 3:1 2:1 1:1 1:2 1:3 PPL($\downarrow$) 5.89 5.86 6.07 6.73 7.45 8.62
Table 6: The perplexity of different $\lambda$ on LLaMA2-7B.
$\lambda$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $5\times 10^{-3}$ $10^{-2}$ $5\times 10^{-2}$ $10^{-1}$ PPL($\downarrow$) 7.47 6.04 5.86 5.93 6.12 7.26 10.55
Data Ratio of Regualrization and RFT With all other settings consistent with Section 4.2, we only varied the data ratio, and the results are shown in Table 5. It can be observed that when the ratio is 3:1, the model achieves the lowest perplexity. When the ratio is 4:1, the increase in perplexity is less pronounced, but when the ratio is 1:1, the perplexity increases significantly by 15%. These results underscore the importance of the regularization process relative to the RFT applied after pruning.
Trade Off Between Language Modeling and Regularization Keeping all other settings consistent with Section 4.2, we evaluate the model’s performance by varying $\lambda\in[10^{-5},10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}]$. As shown in Table 6, when $\lambda=10^{-5}$, the regularization loss is too small to effectively shift important values, leading to a significant increase in perplexity after pruning. In contrast, when $\lambda=10^{-1}$, excessive regularization compromises the model’s language modeling ability, causing a notable performance decline post-pruning. Optimal performance is achieved at $\lambda=10^{-3}$, demonstrating the importance of balancing language modeling loss and regularization loss for effective pruning.
5 Conclusion
In this paper, we introduce a novel pruning paradigm: first applying regularization to the pruned components, followed by pruning and then RFT. Unlike traditional paradigms that first prune based on importance and then apply RFT, DReSS transfers critical information from the pruned parameter space to the remaining components during the regularization, effectively mitigating the irreversible performance degradation caused by information loss. DReSS demonstrates superior performance in both generation and zero-shot tasks, significantly outperforming existing pruning methods. For instance, DReSS surpasses the powerful LLM Surgeon by 20% in perplexity on LLaMA2-7B. On OPT-13B, with a pruning ratio of 25%, the average accuracy drops by only 1%, while achieving a 1.41× speedup compared to the dense model. Moreover, DReSS requires only 25% of the data to achieve comparable performance to previous methods, substantially reducing both data and computational costs while minimizing the reliance on RFT.
Impact Statement
Pruning plays a crucial role in model compression. DReSS is an efficient and low-overhead pruning method, which enables smaller models to achieve competitive performance, especially in resource-constrained scenarios and it also provides some insights for the development of LLMs.